20m.rne07 is reporting non-zero diagnostic flags since this afternoon and thus all data are being flagged as bad. pio didn't help.
P.rsw02 is reported as bad, but data look good logging into the dsm. Not sure what configuration isn't working.
Qsoil.tnw09 still negative.
Several TRHs out (I'll update the list later)